r - s + 2t = -1
2r + 2s - t = 0
r + s = 5
2r + 2t = 4

Answer by Alan3354(69443)

r - s + 2t = -1
2r + 2s - t = 0
r + s = 5
2r + 2t = 4
===============
You have more eqns than variables.